OpenStudy (anonymous):

evaluate cos(tan^-1(-1)

OpenStudy (tkhunny):

Always draw a right triangle. From the drawing, and considering the Quadrant and Reference Angle, we see that: \(\tan^{-1}(-1) = -\dfrac{\pi}{4}\) Again, considering the Quadrant and Reference Angle, we see that: \(\cos(-\pi/4) = \dfrac{1}{\sqrt{2}}\)